solve the second order differential equation
d
2
y
d
x
2

4
y
=
12
x
,
y
(
0
)
=
4
,
y

(
0
)
=
1
d^2y/dx^2−4y=12x,y(0)=4,y′(0)=1
1
Expert's answer
2019-04-17T03:50:30-0400

The general solution of this nonhomogeneous differential equation can be written as

"\text{y}\left( x \right)={{y}_{c}}\left( x \right)+{{y}_{p}}\left( x \right)"

where yc is the general solution of the complementary equation

"\frac{{{d}^{2}}y}{d{{x}^{2}}}-4y=0"

and  yp is a particular solution of equation

"\frac{{{d}^{2}}y}{d{{x}^{2}}}-4y=12x"

First find the general solution of the complementary equation. For this we solve the auxiliary equation

"{{r}^{2}}-4=0\,\,\,=>\,\,{{r}_{1}}=2,\,\,\,{{r}_{2}}=-2"

So, the solution of the complementary equation is

"{{y}_{c}}={{c}_{1}}{{e}^{2x}}+{{c}_{2}}{{e}^{-2x}}"

Since the right side of the original equation is a polynomial of degree 1, i.e. 12x, we seek a particular solution of the form

"{{y}_{p}}\left( x \right)=Ax+B"

Then

"{{{y}'}_{p}}\left( x \right)=A,\,\,\,{{{y}''}_{p}}\left( x \right)=0"

So, substituting into the given differential equation, we have

"0-4\left( Ax+B \right)=12x"

or

"-4Ax-4B=12x"

Polynomials are equal when their coefficients are equal. Thus

"A=-3,\,\,\,B=0"

Therefore, a particular solution is

"{{y}_{p}}\left( x \right)=-3x"

and the general solution is

"y\left( x \right)={{c}_{1}}{{e}^{2x}}+{{c}_{2}}{{e}^{-2x}}-3x"

Now we find c1 and c2. Use the condition y(0)=4

"y\left( 0 \right)={{c}_{1}}{{e}^{2\cdot 0}}+{{c}_{2}}{{e}^{-2\cdot 0}}-3\cdot 0={{c}_{1}}+{{c}_{2}}=4"

Find y′(x) and then use the condition y′(0)=1

"{y}'\left( x \right)={{\left( {{c}_{1}}{{e}^{2x}}+{{c}_{2}}{{e}^{-2x}}-3x \right)}^{\prime }}=2{{c}_{1}}{{e}^{2x}}-2{{c}_{2}}{{e}^{-2x}}-3"

"{y}'\left( 0 \right)=2{{c}_{1}}{{e}^{2\cdot 0}}-2{{c}_{2}}{{e}^{-2\cdot 0}}-3=2{{c}_{1}}-2{{c}_{2}}-3=1"

From the first equation we get

"{{c}_{2}}=4-{{c}_{1}}"

and substitute it into the second equation

"2{{c}_{1}}-2\left( 4-{{c}_{1}} \right)-3=1"

or

"4{{c}_{1}}=12"

Then

"{{c}_{1}}=3,\,\,\,{{c}_{2}}=4-3=1"

So, the solution of the given equation is

"y\left( x \right)=3{{e}^{2x}}+{{e}^{-2x}}-3x"